The weights of a large number of miniature poodle are approximately normally distributed with a mean of 8 kilograms and a standard deviation of 0.9 kilograms. If measurements are recorded to the nearest tenth of a kilogram, find the fraction of these poodle

1
Expert's answer
2022-04-29T10:01:54-0400

Let "X="the weight of a miniature poodle: "\sim N(\mu, \sigma^2)."  

a.

"P(X>9.5)=1-P(Z\le\dfrac{9.5-8}{0.9})"

"\approx0.0478"

b.

"P(X\le8.6)=P(Z\le\dfrac{8.6-8}{0.9})\approx0.7475"

c.


"P(7.3\le X\le9.1)=P(Z\le\dfrac{9.1-8}{0.9})"

"-P(Z\le\dfrac{7.3-8}{0.9})\approx0.6708"